Tell whether y=-3x^2+18x-20 has a minimum value or a maximum value?Then find the minimum or maximum value.
Answer by lwsshak3(11628)

Tell whether y=-3x^2+18x-20 has a minimum value or a maximum value?Then find the minimum or maximum value.
**
Second degree equations are parabolas.
Standard form of equation for a parabola which opens downwards: y=-A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex. A is a multiplier which affects the slope of the curve. The negative lead coefficient means the parabola opens downward and therefore, has a maximum value.
y=-3x^2+18x-20
complete the square
y=-3(x^2-6x+9)-20+27
y=-3(x-3)^2+7
..
given parabola has a maximum=7 at x=3